find the equation of a line passing through (2,-3) (6, 5)
Answer by stanbon(75887)

find the equation of a line passing through (2,-3) (6, 5)
----------------------
slope = (5--3)/(6-2) = 8/4 = 2
----
Form: y = mx + b
Solve for "b":
5 = 2*6 + b
b = -7
-----
Equation:
y = 2x-7
===================
Cheers,
Stan H.